OpenStudy (anonymous):

A wire is first bent into the shape of a triangle. Each side of the triangle is 8 cm long. Then the wire is unbent and reshaped into a rectangle. If the length of the rectangle is 8 cm , what is its width?

OpenStudy (anonymous):

@iheartfood

OpenStudy (***[isuru]***):

fisrt find the perimeter of the triangle... iit will give u the length of the wire..  since the perimeter is equal to the sum length of all sides... the perimeter of the triangle will be (8 + 8 + 8 ) = 24 cm which means that the length of the wire is 24 cm.. now when it shaped in to a rectangle... the perimeter of the rectangle = the length of the wire... = 24 cm and the slandered way of finding the perimeter of a rectangle is (2 x width ) + ( 2 x length ) if length = l and width = w then perimeter = 2l + 2w since we know length = l = 8 then perimeter = (2x8) + 2w 24 = 16 + 2w 2w = 8 w = 4 cm = width got it :) ?

OpenStudy (anonymous):

yup thanks

OpenStudy (***[isuru]***):

u r welcome!